# PhysioWave: A Multi-Scale Wavelet-Transformer for Physiological Signal Representation

### v2

@@ E.1 Visualization of sEMG Pretraining
The following visualizations demonstrate the pretraining process for EMG signals, where the model learns to reconstruct the masked regions of the input signal, specifically focusing on the restoration of multi-scale wavelet decomposition features, $\mathrm{Spec}(X)$. The images show different stages of the reconstruction task, including the target signal, the model’s prediction, the masked areas, and the final reconstructed signal. These visualizations illustrate the model’s ability to restore missing information from wavelet-decomposed features.
@@
Figure  7 shows the results for the 10th training epoch and the 1000th batch of the EMG pretraining process. As seen in the image, the model has learned to effectively restore the missing parts of the signal, showcasing its ability to handle the wavelet decomposition and its corresponding multi-scale features.
@@
Figure  8 displays the results for the 20th training epoch and the 1000th batch of the EMG pretraining process. At this stage, the model shows improved performance in reconstructing the masked regions, indicating its enhanced understanding of the signal’s multi-scale features and its ability to restore more complex signal components.
These visualizations highlight the progression of the model’s learning ability during the pretraining process. The first figure  7 demonstrates the model’s early-stage capability to restore missing multi-scale features, while the second figure  8 shows the improvements made by the model as it continues training. This process emphasizes the model’s capacity to learn to restore wavelet-decomposed features, which is critical for downstream applications like EMG signal analysis and classification.
E.2 Visualization of EPN612 Experimental Results
@@
Figure 9 shows the test confusion matrix for the EPN612 dataset. Per-class recalls remain consistently high ($\geq$87.4%, typically $>$92%), with the highest recall achieved for noGesture (99.4%). Most misclassifications occur between semantically similar dynamic gestures: (a) open is occasionally confused with waveOut and pinch, and (b) pinch is sometimes predicted as open. The sparse distribution of off-diagonal entries indicates that the model has successfully learned discriminative representations for the remaining gesture classes.
@@
Figures 12, 13, and 14 present comprehensive 3D visualizations of the learned feature embeddings using three different dimensionality reduction techniques: Principal Component Analysis (PCA), t-Distributed Stochastic Neighbor Embedding (t-SNE), and Uniform Manifold Approximation and Projection (UMAP). Each visualization displays the same feature space from four distinct viewing angles (elevation/azimuth combinations of 20°/45°, 45°/90°, 60°/135°, and 30°/225°) to provide complete spatial understanding of the clustering patterns.
@@
The t-SNE visualization (Figure 13) emphasizes local neighborhood relationships and reveals non-linear clustering structures. The method successfully preserves local similarities, creating tight, well-separated clusters for each gesture class. Notably, the noGesture class forms a dense, spherical cluster at the center, while dynamic gestures are arranged in distinct regions around the periphery, suggesting strong discriminative power in the learned representations.
@@
The UMAP visualization (Figure 14) balances both local and global structure preservation, revealing the topological organization of the feature space. UMAP shows excellent class separation with noGesture forming a highly concentrated cluster, and dynamic gesture classes organized in a manifold-like structure that preserves both local neighborhoods and global relationships between semantically related gestures.
@@
Across all three visualization methods, several consistent patterns emerge: (1) the noGesture class demonstrates exceptional separability from active gestures, (2) dynamic gesture classes maintain distinct clustering while showing logical spatial relationships based on gesture similarity, and (3) the learned feature representations exhibit strong discriminative properties that facilitate effective classification. The multi-view presentation confirms the robustness of these clustering patterns across different spatial perspectives.
Appendix F LIMITATIONS
@@
While the proposed PhysioWave architecture represents a significant advancement in the processing of physiological signals, it is not without its limitations. Our current approach focuses on large-scale models specifically trained for electromyography (EMG) and electrocardiography (ECG) signals, both of which have demonstrated state-of-the-art performance across several downstream tasks. However, this approach is currently limited to these two modalities, and other important physiological signals, such as electroencephalography (EEG) and photoplethysmography (PPG), remain underexplored.
@@
Another key limitation lies in the fact that our work primarily focuses on separate models for different physiological signals. Although our unified framework integrates pre-trained EMG and ECG models with an existing EEG encoder, the design is not fully generalized across all biosignal modalities. This calls for further exploration into the development of a universal, multi-modal model capable of seamlessly processing a wide variety of physiological signals. A model of this nature could function as a comprehensive "physiological diagnostic" system, akin to a highly specialized medical expert, diagnosing and interpreting a broad spectrum of physiological signals across different conditions. Developing such a universal model would not only enhance the robustness of signal interpretation but also streamline the diagnostic process across different sensor modalities.
@@
We believe that addressing these gaps by expanding the range of modalities incorporated into our framework and further investigating the potential for a unified, multi-modal biosignal model will significantly improve the versatility and applicability of our approach in real-world clinical and health monitoring settings. Future research could focus on training models for a wider array of biosignals, as well as enhancing the multi-modal integration capabilities to create a truly universal model for physiological signal processing.

### v4

@@ F.2 Visualization of sEMG Pretraining
The following visualizations demonstrate the pretraining process for EMG signals, where the model learns to reconstruct the masked regions of the input signal, specifically focusing on the restoration of multiscale wavelet decomposition features, $\mathrm{Spec}(X)$. The images show different stages of the reconstruction task, including the target signal, the prediction of the model, the masked areas, and the final reconstructed signal. These visualizations illustrate the model’s ability to restore missing information from wavelet-decomposed features[11].
@@
Figure  9 shows the results for the 10th training epoch and the 1000th batch of the EMG pretraining process. As seen in the image, the model has learned to effectively restore the missing parts of the signal, showcasing its ability to handle the wavelet decomposition and its corresponding multiscale features.
@@
Figure  10 displays the results for the 20th training epoch and the 1000th batch of the EMG pretraining process. At this stage, the model shows improved performance in reconstructing the masked regions, indicating its enhanced understanding of the signal’s multiscale features and its ability to restore more complex signal components.
These visualizations highlight the progression of the model’s learning ability during the pretraining process. The first figure  9 demonstrates the model’s early-stage capability to restore missing multiscale features, while the second figure  10 shows the improvements made by the model as it continues training. This process emphasizes the model’s ability to learn to restore wavelet-decomposed features, which is critical for downstream applications such as EMG signal analysis and classification.
F.3 Visualization of EPN612 Experimental Results
@@
Figure 11 shows the test confusion matrix for the EPN612 dataset. Per-class recalls remain consistently high ($\geq$87.4%, typically >>92%), with the highest recall achieved for noGesture (99.4%). Most misclassifications occur between semantically similar dynamic gestures: (a) open is occasionally confused with waveOut and pinch, and (b) pinch is sometimes predicted as open. The sparse distribution of off-diagonal entries indicates that the model has successfully learned discriminative representations for the remaining gesture classes.
@@
Figures 14, 15, and 16 present comprehensive 3D visualizations of the learned feature embeddings using three different dimensionality reduction techniques: Principal component analysis (PCA), t-distributed stochastic neighboring incorporation (t-SNE) and uniform manifold approximation and projection (UMAP). Each visualization shows the same feature space from four different viewing angles (elevation / azimuth combinations of 20/45, 45/90, 60/135, and 30/225) to provide complete spatial understanding of the clustering patterns.
@@
The visualization of t-SNE (Figure 15) emphasizes local neighborhood relationships and reveals non-linear clustering structures. The method successfully preserves local similarities, creating tight, well-separated clusters for each gesture class. In particular, the noGesture class forms a dense spherical cluster in the center, while dynamic gestures are arranged in distinct regions around the periphery, suggesting strong discriminative power in the learned representations.
@@
The visualization of UMAP (Figure 16) balances the preservation of local and global structures, revealing the topological organization of the feature space. UMAP shows excellent class separation with noGesture forming a highly concentrated cluster, and dynamic gesture classes organized in a manifold-like structure that preserves both local neighborhoods and global relationships between semantically related gestures.
@@
Across the three visualization methods, several consistent patterns emerge: (1) the noGesture class demonstrates exceptional separability from active gestures, (2) dynamic gesture classes maintain distinct clustering while showing logical spatial relationships based on gesture similarity, and (3) the learned feature representations exhibit strong discriminative properties that facilitate effective classification. The multiview presentation confirms the robustness of these clustering patterns across different spatial perspectives.
F.4 Model Interpretability via GradCAM Analysis
To ensure clinical interpretability and validate that our model learns physiologically meaningful patterns, we employed Gradient-weighted Class Activation Mapping (GradCAM) [61] to visualize the temporal regions most influential to classification decisions. This analysis provides crucial insights into whether the model attends to clinically relevant ECG segments corresponding to established diagnostic criteria.
F.4.1 Methodology
We applied GradCAM to the final transformer block of our trained PhysioWave-ECG model, generating importance heatmaps across the temporal dimension for representative samples from the Georgia ECG dataset. The analysis focused on cardiac rhythm classes with distinct temporal signatures, including Normal Sinus Rhythm (NSR), T-wave Inversion (TInv), and ST-segment Depression (STD). For each sample, we computed gradient-based importance scores and overlaid them on the original ECG signals to identify regions of maximum model attention.
F.4.2 Temporal Attention Patterns
Figure 17 illustrates representative GradCAM visualizations for different cardiac conditions, revealing distinct attention patterns that align with clinical diagnostic features:
(a) Normal Sinus Rhythm (NSR)
(b) T-wave Inversion (TInv)
(c) ST-segment Depression (STD)
(d) Complex Arrhythmia Pattern
Figure 17: GradCAM visualization results showing model attention patterns for different cardiac conditions. Red regions indicate high importance scores, demonstrating that the model focuses on clinically relevant temporal segments for each pathology.
F.4.3 Clinical Validation of Learned Features
The GradCAM analysis revealed distinct attention patterns that strongly align with established clinical diagnostic criteria across different rhythm classes. For T-wave Inversion (TInv), the model’s attention was predominantly concentrated in the 0–400ms window, precisely targeting the early signal segments where T-wave morphology deviates from normal patterns. This focused attention confirms that the model successfully identifies repolarization abnormalities, a critical feature for diagnosing T-wave pathologies[62].
In contrast, Normal Sinus Rhythm (NSR) samples exhibited distributed attention across periodic cardiac cycles with recurring importance spikes at regular intervals. This pattern indicates the model’s reliance on complete P-QRS-T morphology rather than isolated peaks, demonstrating its ability to recognize the harmonious relationship between cardiac components that characterizes healthy rhythm patterns. The periodic nature of the attention maps further validates the model’s understanding of rhythmic regularity[62].
For ST-segment depression (STD), the highest activation appeared consistently in the early temporal regions corresponding to post-QRS intervals, precisely where ST-segment changes manifest in clinical practice. This targeted attention demonstrates the successful isolation of diagnostically critical ST windows of the model, which is essential for detecting ischemic changes[63]. Complex arrhythmias presented variable attention patterns with multiple focal points distributed throughout the temporal sequence, reflecting the model’s adaptive feature extraction capabilities when faced with irregular rhythm patterns and compound pathologies.
F.4.4 Quantitative Analysis of Attention Distribution
To quantify the focus of the model on clinically relevant intervals, we analyzed the temporal distribution of GradCAM activation across 100 samples per class. The analysis revealed:
  • Pathology-specific localization: For ST-segment abnormalities, 78% of peak activations occurred within the 100–300ms post-QRS window, corresponding to the ST-segment interval in standard ECG interpretation.
  • Periodic attention for regular rhythms: NSR samples showed autocorrelation peaks in GradCAM maps at intervals matching the heart rate (mean R-R interval), confirming the model’s recognition of rhythmic patterns.
  • Early temporal focus for repolarization disorders: T-wave abnormalities triggered concentrated attention in the first 40% of the signal window, aligning with the expected temporal location of T-wave morphology.
F.4.5 Clinical Implications
The GradCAM analysis provides three key insights for clinical deployment:
  1. 1.
    Diagnostic Transparency: The visualization enables cardiologists to verify that automated diagnoses are based on appropriate ECG segments rather than spurious correlations or artifacts.
  2. 2.
    Trust Building: By demonstrating attention to the same temporal landmarks used in manual interpretation (P-waves, QRS complexes, ST-segments, T-waves), the model establishes credibility with clinical practitioners.
  3. 3.
    Educational Value: The attention maps can serve as teaching tools, highlighting critical diagnostic regions for medical students and residents learning ECG interpretation.
These results confirm that our wavelet-based transformer architecture not only achieves high classification accuracy but also learns interpretable, physiologically grounded features. The alignment between algorithmic attention patterns and established clinical knowledge validates the model’s suitability for computer-aided diagnosis in real-world healthcare settings, where interpretability is paramount for adoption and regulatory approval.
Appendix G LIMITATIONS
@@
While the proposed PhysioWave architecture represents a significant advance in the processing of physiological signals, it is not without limitations. Our current approach focuses on large-scale models specifically trained for electromyography (EMG) and electrocardiography (ECG) signals, both of which have demonstrated state-of-the-art performance in several downstream tasks. However, this approach is currently limited to these two modalities, and other important physiological signals, such as electroencephalography (EEG) and photoplethysmography (PPG), remain underexplored.
@@
Another key limitation lies in the fact that our work focuses primarily on separate models for different physiological signals. Although our unified framework integrates pre-trained EMG and ECG models with an existing EEG encoder, the design is not fully generalized across all biosignal modalities. This calls for further exploration into the development of a universal, multimodal model capable of seamlessly processing a wide variety of physiological signals. A model of this nature could function as a comprehensive "physiological diagnostic" system, akin to a highly specialized medical expert, diagnosing and interpreting a broad spectrum of physiological signals across different conditions. Developing such a universal model would not only enhance the robustness of signal interpretation but would also streamline the diagnostic process across different sensor modalities.
@@
We believe that addressing these gaps by expanding the range of modalities incorporated into our framework and further investigating the potential for a unified, multimodal biosignal model will significantly improve the versatility and applicability of our approach in real-world clinical and health monitoring settings. Future research could focus on training models for a wider array of biosignals, as well as enhancing the multi-modal integration capabilities to create a truly universal model for physiological signal processing.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We point out the limitations of the work and explain why they are important to consider in Appendix  G.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: This paper has detailed instructions on how to reproduce the main experimental results in Appendix  C and supplementary materials.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We release all code and data preprocess scripts in the supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper specify all the training and test details in Appendix  E
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The results are accompanied by error bars, confidence intervals, or statistical significance tests.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper provide sufficient information on the computer resources in Appendix  C
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper discusses the potential positive societal impacts of the proposed system, particularly in health monitoring, clinical diagnostics, and personalized medicine. The use of multi-modal biosignal processing and self-supervised learning frameworks such as PhysioWave can significantly enhance diagnostic accuracy, leading to improved healthcare outcomes, more personalized treatments, and broader accessibility to advanced health technologies.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the original paper of existing assets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release our code and paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.